Problem 1: The 60 deg strain rosette is mounted on a steel plate (E = 30e6 psi, G = 11.5e6 psi). Readings obtained from eachits 45 deg rosette.Not 60 degree
Answer #1

Solution in Criven that in {=737*156 E = 30x106psi G= 11.5*108 Rsi & Ex= 389x16 450 450 &z=-2904156 &,=Ex Aug 389*166 Ex=7374E = Ex = 389*156 E = -2904106 o -45° we have E = Exloso + + Ey Sinke & Vay Coso. Sino -29000 6 3899100 cost (45) & Ey sin 26-ܓܢ Gm 6 E дож to® » 229 м іѕ6 Р. 16x = 11670 psi Ang S, 6 = €€ = 62ж 66x3ож “ by = 1740 psi Any — 6 ||:5% 10% 1025 ж * 166 у